# Quartic soliton solutions of a normal dispersion based mode-locked laser

### v1

@@ I Introduction
Quartic solitons are solitons of optical models whose dispersion is dominated by a fourth order term, whereas the term pure quartic solitons is reserved for models where only the fourth order dispersion is considered. They have been predicted [1, 2, 3] and observed [4, 5] in conservative models but they have been more acclaimed in dissipative models for their advantages in mode-locked lasers and Kerr soliton frequency combs whenever the fourth order dispersion (4OD) is negative (anomalous 4OD). In fact, in mode-locked lasers the negative quartic dispersion may give rise to pulses whose energy scales inversely with the width cubed, enabling very energetic ultrashort pulses [6, 7, 8]. For Kerr combs presenting negative 4OD, the same energy-width relation was found but more importantly solitons spectra is flatter [9]. Quartic solitons have also been studied in more complex models like vectorial models for mode-locked lasers [10] and a fiber resonator model allowing orbital angular momentum carrying modes [11]. In case of positive 4OD (also named normal 4OD), solitons do not exist for conservative models based on Kerr nonlinearity [12] but do exist in dissipative models. A triangular-shaped pulse with a double peak spectrum has been reported by Runge et al. [13] for a modified nonlinear Schrödinger equation (NLSE) comprising positive 4OD, Kerr effect and gain as well as for a mode-locked laser lumped model. Note that, this kind of pulses was already reported as non-soliton solutions of the NLSE with positive second order dispersion (2OD) and positive 4OD [14]. Another type of soliton solution for mode-locked laser models with positive 4OD was reported in [15, 7] which also presents a double-peaked spectrum but whose temporal profile consists of a sech central part on top of a large pedestal. Recently and while we were obtaining the results here reported, existence regions for both solutions by considering the complex Ginzburg-Landau equation (CGLE) were reported elsewhere [16]. In [15], a third kind of solution that is asymmetrical was reported. Nevertheless, all the reported solutions of mode-locked lasers dominated by normal 4OD were not yet sufficiently studied in terms of pulse characteristics, energy-width relation, regions of existence and stability, coexistence and onset from continuous wave evolution.
@@
Here, we report a thorough study of soliton solutions of a distributed model for mode-locked lasers [17, 7] dominated by normal fourth order dispersion. Two of the solutions are similar to the above referred solutions, namely, the triangular-shaped and the sech-pulse above a pedestal. We have found that they coexist in some region of the parameter space, at which they are also stable, manifesting dynamics of bistability. The other ones were obtained by solving the associated ordinary differential equation but are shown to be unstable. Three main branches of those solutions (one always unstable and two stable in some regions) are characterized in terms of amplitude, phase and spectrum profiles and in terms of energy exchange with the exterior and within themselves. The other branches occur on a bifurcation region of parameters and only the amplitude profiles are presented. The parameter regions of existence and stability of the two stable solutions are determined and their energy-width scaling is analyzed. The model is also studied concerning the modulational instability of its continuous wave (cw) solution and the relation of the modulational instability and the onset of solitons is analyzed. The influence of lower order dispersion terms, namely, second order dispersion (2OD) and third order dispersion (3OD), on the existence and stability of the stable solitons is investigated.
@@ II Model equations and methods
Our search for pulse solutions has used two different methods. Full integration of the Eq. (3) using a localized input and integration of the ordinary differential equation (ODE) that resulted from Eq. (3) by applying a similarity variable transformation, namely $q(Z,T)=F(T)e^{i\sigma Z}$, where $F$ is a complex amplitude and $\sigma$ a real propagation constant. The resulting ODE is given by
@@
The PDE in Eq. (3) was solved through pseudo-spectral methods and the ODE in Eq. (5) was integrated using the Newton conjugate-gradient (NCG) method developed by J. Yang [18, 19], starting from hyperbolic secant profiles or from soliton solutions obtained in previous simulations with similar parameters.
@@ III.1 Types of Solutions
The application of the methods referred in the previous section allowed us to find different types of soliton solutions of the model in Eq. (1) with null 2OD and 3OD and normal 4OD. Three main solution branches are studied more carefully and are named according to their relative peak amplitude as low amplitude soliton (LAS), medium amplitude soliton (MAS) and high amplitude soliton (HAS). We have kept fixed $d_{\text{SA}}=0.3$, $\gamma=0.005$ W-1m-1, $L=1$ m and $k_{\text{OC}}=-\text{ln}(0.3)$ as well as the saturation power $P_{\text{sat}}=80$ W and the 4OD coefficient $\beta_{4}=0.08$ ps4m-1. This choice of parameters was based on previous works such as [17] where the distributed model (with 2OD only) was introduced and [15] which reported normal quartic solitons in mode-locked lasers.
The power profile given by $|W(t)|^{2}$, chirp and spectrum profiles are shown in graphs of Figs. 1, 2 and 3. The LAS power profile has wider tails than a sech2, linear chirp at the pulse peak and constant chirp at the tails and its spectral profile shows one single peak at the carrier frequency (Fig. 1). As mentioned in the introduction, all the solutions of this type were found to be unstable, having one unstable eigenvalue that is purely imaginary.
@@
The solutions named MAS have different characteristics depending on the value of $T_{2}$. For lower $T_{2}$, the power profile is close to a sech2 ($P_{p}\,\text{sech}^{2}(t)$, with $P_{p}$ being the peak power) on top of a pedestal. The chirp is also linear at the pulse peak position and constant at the tails (similar to the LAS chirp profile) and the spectrum exhibits two distinguished peaks as reported before [15] (Figs. 2 and 2). For higher $T_{2}$, the power profile is close to $P_{p}\,\text{sech}^{2}(t/\zeta)$, with $\zeta>1$, thus, it is wider than the $P_{p}\,\text{sech}^{2}(t)$ that adjusts to the central part of the MAS solutions for lower $T_{2}$. The chirp is also similar to the ones referred before for LAS and MAS with low $T_{2}$ but showing less structure, as may be observed in Fig. 2, and the spectrum is single peaked (Fig. 2). These characteristics were already reported by us in [7] and are analogous to those previously found in [15, 16].
@@
$W(t)=a\exp\left[-\left(\frac{|t|}{c}\right)^{b}\right],$ (10)
@@
Figure 4: Values of $a$, $b$ and $c$ obtained from fitting Eq. (10) to HAS amplitude profiles as function of $T_{2}$. The values of $g_{0}$ (m-1) are indicated in the legends.
@@ III.2 Energy Flow
In order to understand the energy exchange dynamics to the exterior and within the pulse, we have followed [20] and proved the equality
$\frac{\partial s}{\partial Z}+\frac{\partial j}{\partial T}=P,$ (11)
where $s=|q|^{2}$ is a density of energy, which is constant along $Z$ for stationary solitons, $j$ is an internal flux of energy given by
$j=i\frac{D_{4}}{24}\left(qq_{TTT}^{*}-q_{TTT}q^{*}+q_{TT}q_{T}^{*}-q_{T}q_{TT}^{*}\right)$ (12)
and $P$ is the density of energy that enters the system from the exterior given by
@@
The graphs of Fig. 5 show $j$ and $P$ for all the solutions of Figs. 1-3. For all solutions, the energy enters through the center of the pulse and then is routed to the tails where it is dissipated. This behaviour was expected since the saturable absorber is saturated in the center, allowing for nonlinear gain to manifest (thus, energy is entering the pulse at the center), and is unsaturated at the tails, where losses will occur.
Figure 5: Density of energy generation (blue line) and internal flux of energy (red line) both graphed versus temporal position of the LAS for $g_{0}=1.465$ m-1, $T_{2}=100$ fs 5, MAS for $g_{0}=1.465$ m-1, $T_{2}=100$ fs 5, MAS for $g_{0}=1.485$ m-1, $T_{2}=550$ fs 5 and HAS for $g_{0}=1.465$ m-1, $T_{2}=100$ fs 5. In all, cases $P_{\text{sat}}=80$ W. All the quantities are dimensionless. The amplitude profiles are shown without scale.
@@ III.3 Existence, stability and bistability
The three main solutions are related by a bifurcation kind of dynamics that is shown in Fig. 6 for the particular value of $g_{0}=1.465~{}\text{m}^{-1}$. The graph shows the peak power as $T_{2}$ is swept for the three types of solutions. The lower branch corresponds to LAS which is connected to the MAS branch at a higher threshold $T_{2}$ value (approximately 200 fs in the case shown in the graph). There are stable and unstable MAS solutions, stable for higher $T_{2}$ but unstable below a particular value of $T_{2}$ (approximately 100 fs in the case shown in the graph). The HAS bifurcates in a complicated way from the branch of MAS. LAS, MAS and HAS solutions can coexist, either all three at once, or in pairs, as well as with the other more restricted branches of solutions represented in Fig. 6. The solutions in these other branches are all unstable and were found by integrating the ODE for different values of $T_{2}$, starting the NCG method with a previous solution, and modifying $T_{2}$ in different steps and in opposite directions. The profiles of the pulses in those branches are shown in Fig. 6 and all of them exhibit the sharp pulse on top of a larger part that is characteristic of the HAS solutions. Solutions identified as C and D are actually quite similar to the HAS solutions. In the other solutions represented, the larger part of the pulse is seen to be composed of two different regions. This characteristic is more pronounced in pulse F.
@@
MAS and HAS are both stable in a region of parameters, as shown below, producing bistability behaviors. One of these behaviors is hysteresis as may be observed in the results shown on Fig. 6. We solved Eq. (3) using inputs from previous simulations for close values of $T_{2}$. The results labeled as ’Increasing $T_{2}$’ correspond to the output solutions obtained with inputs of lower $T_{2}$ and the ones labeled as ’Decreasing $T_{2}$’ correspond to the output solutions obtained with inputs of higher $T_{2}$. In fact, in the coexistence region the results obtained with different inputs do not coincide. In the direction of decreasing $T_{2}$, the jump between the two type of solutions is abrupt. However, in the other direction, there is a range of $T_{2}$ (105 to 109 fs) for which the simulation did not produce a stationary pulse but instead a pulse that evolved in $Z$ with peak and form variations. This behavior can be explained by the existence of different branches of unstable solutions in that region.
@@
Figure 6: 6 Bifurcation diagram showing the peak power versus $T_{2}$ with several branches of solutions (including LAS, MAS and HAS). 6 Power profiles for the solutions identified in 6. 6 Hysteresis shown by the peak value of the output pulses obtained by solving Eq. (3) with inputs from higher $T_{2}$ (Decreasing $T_{2}$) and inputs from lower $T_{2}$ (Increasing $T_{2}$). The results shown correspond to $P_{\text{sat}}=80$ W and $g_{0}=1.465$ m-1.
@@
Restricting our attention to possible stable solutions, namely, to MAS and HAS solutions, we scanned the $(g_{0},T_{2})$ space to find regions at which those solutions exist and are stable. The regions are those in Figs. 7 and 8 where energy and temporal width are also shown. For increased clarity, the boundaries of the graphs should be explained. The right boundary at $g_{0}=1.50~{}\text{m}^{-1}$ is an actual boundary of existence for both types of solitons since, for higher $g_{0}$, the linear loss would turn to linear gain and the background would become unstable, and even if solitons exist, they would not survive in the unstable background. The upper boundary of MAS, for $T_{2}$ higher than 880 fs and $g_{0}>1.48$, is not an actual boundary but only a limit at which the scanning of the $(g_{0},T_{2})$ space was stopped.
Figure 7: Contour plots of the energy in pJ 7 and pulse width in ps 7 for MAS.
@@
Concerning the region of existence and stability of MAS, it was found to be composed by two distinguishable bands. The first, exists for a $g_{0}$ range between 1.443 and 1.475 m-1, being very limited to $T_{2}$ values around 100 fs. The second band exists in a much broader $T_{2}$ region, with $T_{2}\gtrsim 100$, but is much more limited in $g_{0}$, ranging from 1.475 to 1.5 m-1. Note that the results for $\beta_{4}>0$ presented in [7] were for MAS on the first referred thin band of existence. The lower boundary of the MAS region corresponds to a transition to unstable solutions at which two complex conjugated stability eigenvalues cross the real axis from the stable half plane to the unstable half plane. On the other hand, above the upper boundary, no soliton solutions were found. Figs. 7 and 7 also have contour lines that show that, for lower values of $T_{2}$ the most energetic pulses are also the shortest ones.
@@
Figure 8: Contour plots of base 10 logarithm of the energy in pJ 8 and pulse width in ps 8 for the HAS.
@@
As we referred above, the overlap of regions of stable MAS and LAS is not null, i.e., there are parameter values for which both MAS and LAS solutions exist and are stable. This parameter region of bistability is better observed in Fig. 9, showing it to be small when compared with regions of existence of each separate type of soliton.
@@
Figure 10: Evolution from an LAS solution to a HAS solution (passing through an MAS solution) for $g_{0}=1.47~{}\text{m}^{-1}$ and $T_{2}=100~{}\text{fs}$. 10 Evolution of the peak power in $z$ and 10 power profiles of the three solutions.

### v2

@@ I Introduction
Quartic solitons are solitons of optical models whose dispersion is dominated by a fourth order term, whereas the term pure quartic solitons is reserved for models where only the fourth order dispersion is considered. They have been predicted [1, 2, 3] and observed [4, 5] in conservative models but they have been more acclaimed in dissipative models for their advantages in mode-locked lasers and Kerr soliton frequency combs whenever the fourth order dispersion (4OD) is negative (anomalous 4OD). In fact, in mode-locked lasers, the negative quartic dispersion may give rise to pulses whose energy scales inversely with the width cubed, enabling highly energetic ultrashort pulses [6, 7, 8]. For Kerr combs presenting negative 4OD, the same energy-width relation was found but more importantly, the soliton spectra are flatter [9]. Quartic solitons have also been studied in more complex models like vectorial models for mode-locked lasers [10] and a fiber resonator model allowing orbital angular momentum carrying modes [11]. In case of positive 4OD (also named normal 4OD), solitons do not exist for conservative models based on Kerr nonlinearity [12] but do exist in dissipative models. A triangular-shaped pulse with a double peak spectrum has been reported by Runge et al. [13] for a modified nonlinear Schrödinger equation (NLSE) comprising positive 4OD, Kerr effect and gain as well as for a mode-locked laser lumped model. Note that, this kind of pulses was already reported as non-soliton solutions of the NLSE with positive second order dispersion (2OD) and positive 4OD [14]. Another type of soliton solution for mode-locked laser models with positive 4OD was reported in [15, 7] which also presents a double-peaked spectrum but whose temporal profile consists of a sech central part on top of a large pedestal. Recently, and while we were obtaining the results here reported, the complex Ginzburg-Landau equation (CGLE) was used to obtain existence regions for the solutions referred above [16] and simulation and experimental results reported the existence of pulses in a fiber laser dominated by positive 4OD [17]. In [15], a third kind of solution, that is asymmetrical, was reported. Nevertheless, all the reported solutions of mode-locked lasers dominated by normal 4OD were not yet sufficiently studied in terms of pulse characteristics, energy-width relation, regions of existence and stability, coexistence and onset from continuous wave evolution.
@@
Here, we report a thorough study of soliton solutions of a distributed model for mode-locked lasers [18, 7] dominated by normal fourth order dispersion. Two of the solutions are similar to the above referred solutions, namely, the triangular-shaped and the sech-pulse above a pedestal. We have found that they coexist in some region of the parameter space, at which they are also stable, manifesting dynamics of bistability. The other ones were obtained by solving the associated ordinary differential equation but are shown to be unstable. Three main branches of those solutions (one always unstable and two stable in some regions) are characterized in terms of amplitude, phase and spectrum profiles and in terms of energy exchange with the exterior and within themselves. The other branches occur on a bifurcation region of parameters and only the amplitude profiles are presented.
The paper is organized as follows. The mathematical and theoretical framework for the distributed model is presented in Section II, alongside a description of the numerical methods used to study the model. Section III characterizes the obtained soliton solutions. Their internal energy flow was studied, parameter regions of existence and stability were found and the energy-width scaling of stable solutions was analyzed. Section IV studies the formation of these solitons through modulational instability. In Section V, the effects of second order dispersion (2OD) and third order dispersion (3OD) on the existence and stability of the stable solitons is investigated. Section VI presents the main conclusions of the work.
@@ II Model equations and methods
Applying the similarity variable transformation given by $q(Z,T)=F(T)e^{i\sigma Z}$, where $F$ is a complex amplitude and $\sigma$ a real propagation constant, we obtain the following ordinary differential equation (ODE)
@@
Our search for pulse solutions has used two different methods: full integration of the Eq. (3) using a localized input and integration of the ODE (5). The PDE in Eq. (3) was solved through pseudo-spectral methods and the ODE in Eq. (5) was integrated using the Newton conjugate-gradient (NCG) method developed by J. Yang [19, 20], starting from hyperbolic secant profiles or from soliton solutions obtained in previous simulations with similar parameters.
@@ III Soliton solutions
This section analyses the quartic soliton solutions of Eq. (1). The first subsection presents pulse profiles and chirp of three types of solitons that were obtained. Subsection III.2 studies the energy exchange dynamics of those solitons. In subsection III.3, regions of existence and stability of those solitons are thoroughly discussed, and behaviors of bistability and hysteris are identified. Other types of soliton solutions that were not thoroughly characterized are also presented. Finally, subsection III.4 studies the energy-width scaling of stable solitons through the variation of equation parameters.
@@ III.1 Types of Solutions
The application of the methods referred in the previous section allowed us to find different types of soliton solutions of the model in Eq. (1) with null 2OD and 3OD and normal 4OD. Three main solution branches are studied more carefully and are named according to their relative peak amplitude as low amplitude soliton (LAS), medium amplitude soliton (MAS) and high amplitude soliton (HAS). We have kept fixed $d_{\text{SA}}=0.3$, $\gamma=0.005$ W-1m-1, $L=1$ m and $k_{\text{OC}}=-\text{ln}(0.3)$ as well as the saturation power $P_{\text{sat}}=80$ W and the 4OD coefficient $\beta_{4}=0.08$ ps4m-1. This choice of parameters was based on previous works such as [18] where the distributed model (with 2OD only) was introduced and [15] which reported normal quartic solitons in mode-locked lasers. In an experimental setting, many of these parameters are intrinsic to the characteristics of the laser itself, be it the setup or its materials. However, effective dispersion management is fundamental for the generation of quartic solitons. The first quartic soliton fiber laser [6] used an intracavity spectral pulse shaper based on a spatial light modulator to make $\beta_{2}$ and $\beta_{3}$ negligible while setting a value of negative $\beta_{4}$. Recently, the same technique was used to generate dissipative quartic solitons in a laser cavity with $\beta_{4}>0$ [17].
The power profile given by $|W(t)|^{2}$, the chirp defined by $-d\phi/dt$ ($\phi$ is the phase of $W(t)$) and the spectrum profiles are shown in graphs of Figs. 1, 2 and 3. The LAS power profile has wider tails than a sech2, linear chirp at the pulse peak and constant chirp at the tails and its spectral profile shows one single peak at the carrier frequency (Fig. 1). As mentioned in the introduction, all the solutions of this type were found to be unstable, having one unstable eigenvalue that is purely imaginary.
@@
The solutions named MAS have different characteristics depending on the value of $T_{2}$. For lower $T_{2}$, the power profile is close to a sech2 ($P_{p}\,\text{sech}^{2}(t)$, with $P_{p}$ being the peak power) on top of a pedestal. The chirp is also linear at the pulse peak position and constant at the tails (similar to the LAS chirp profile) and the spectrum exhibits two distinguished peaks as reported before [15] (Figs. 2 and 2). For higher $T_{2}$, the power profile is close to $P_{p}\,\text{sech}^{2}(t/\zeta)$, with $\zeta>1$, thus, it is wider than the $P_{p}\,\text{sech}^{2}(t)$ that adjusts to the central part of the MAS solutions for lower $T_{2}$. The chirp is also similar to the ones referred before for LAS and MAS with low $T_{2}$ but showing less structure, as may be observed in Fig. 2, and the spectrum is single peaked (Fig. 2). These characteristics were already reported by us in [7] and are analogous to those previously found in [15, 16, 17].
@@
$|W(t)|=a\exp\left[-\left(\frac{|t|}{c}\right)^{b}\right],$ (10)
@@
Figure 4: Values of $a$, $b$ and $c$ obtained from fitting Eq. (10) to HAS amplitude profiles as function of $T_{2}$ and $a^{2}$ versus $c$ for several values of $g_{0}$ (m-1) as indicated in the legends. The different ranges of $T_{2}$ here presented are related to solution existence and stability (see Section C, Fig. 9)
As was already commented in the introduction, the cubic-quintic CGLE has been widely used to model mode-locked lasers. We may reach the cubic-quintic CGLE by expanding the term of the saturable absorber up to the fifth order in $W$, but this should only be valid for small $|W|^{2}/\bar{P}_{\text{sat}}$. The examples shown in Figures 1, 2 and 3 have peak power values of 7.8 W (LAS), 36 and 22 W (MAS) and 140 W (HAS) for a $\bar{P}_{\text{sat}}$ value around 18 W. Thus, among these examples, only the LAS could be reasonable obtained using the cubic-quintic CGLE. Nevertheless, we found LAS with peak values as large as 20 W (Fig. 6), MAS peaking at 50 W, for $g_{0}=1.501~{}\mathrm{m}^{-1}$ and $T_{2}=124$ fs, and HAS peaking at 1600 W, for $g_{0}=1.37~{}\mathrm{m}^{-1}$ and $T_{2}=51$ fs, which are results that invalidate the expansion of the saturable absorber.
@@ III.2 Energy Flow
In order to understand the energy exchange dynamics to the exterior and within the pulse, we use the following continuity relation for equation (3) [21]
$\frac{\partial s}{\partial Z}+\frac{\partial j}{\partial T}=P.$ (11)
In the above equation, $s=|q|^{2}$ is the density of power (power per unit time), which is constant along $Z$ for stationary solitons. $j$ is a flux of density of power, i.e., $j=su$ with $u$ being a velocity (derivative of $T$ in order to $Z$) indicating the direction, in $T$, of the movement of power density as $Z$ is varied, given by
$j=i\frac{D_{4}}{24}\left(qq_{TTT}^{*}-q_{TTT}q^{*}+q_{TT}q_{T}^{*}-q_{T}q_{TT}^{*}\right).$ (12)
The flux of power $j$ is associated with the conservative terms, in this case, the dispersion term, and reveal the direction of the flow of power within the pulse, in parts at which $j$ is positive the flux is in the direction of positive $T$ and if negative the direction of the flux is towards negative $T$. $P$ is the density of power per distance $Z$ that enters or leaves the system from or to the exterior. given by
@@
$P$ is related with the dissipative terms, it is positive in parts of the pulse at which energy is absorbed from the medium and negative in parts at which the pulse dissipates energy.
The graphs of Fig. 5 show $j$ and $P$ for all the solutions of Figs. 1-3. For all solutions, the energy enters through the center of the pulse, in regions of positive $P$ that are identified in the graphs in gray, and then is routed to the tails where it is dissipated. This behaviour was expected since the saturable absorber is saturated in the center, allowing for nonlinear gain to manifest (thus, energy is entering the pulse at the center), and is unsaturated at the tails, where losses will occur.
Figure 5: Density of energy generation (blue line) and internal flux of energy (red line) both graphed versus temporal position of the LAS for $g_{0}=1.465$ m-1, $T_{2}=100$ fs 5, MAS for $g_{0}=1.465$ m-1, $T_{2}=100$ fs 5, MAS for $g_{0}=1.485$ m-1, $T_{2}=550$ fs 5 and HAS for $g_{0}=1.465$ m-1, $T_{2}=100$ fs 5. All the quantities are dimensionless. The dashed lines are the corresponding power profiles with the correct time scale (dimensionless) but with arbitrary units in amplitude, to serve as a reference. The gray dashed zones identify the temporal range at which $P>0$.
@@ III.3 Existence, stability and bistability
The three main solutions are related by a bifurcation kind of dynamics that is shown in Fig. 6 for the particular value of $g_{0}=1.465~{}\text{m}^{-1}$. The graph shows the peak power as $T_{2}$ is swept for the three types of solutions. The lower branch corresponds to LAS which is connected to the MAS branch at a higher threshold $T_{2}$ value (approximately 200 fs in the case shown in the graph). This connection reveals characteristics of a saddle-node bifurcation, at which a stable and an unstable solutions collide at some parameter value and no other solutions exist beyond the same parameter value. There are stable and unstable MAS solutions, stable for higher $T_{2}$ but unstable below a particular value of $T_{2}$ (approximately 100 fs in the case shown in the graph). The HAS bifurcates in a complicated way from the branch of MAS. LAS, MAS and HAS solutions can coexist, either all three at once, or in pairs, as well as with the other more restricted branches of solutions represented in Fig. 6. The solutions in these other branches are all unstable and were found by integrating the ODE for different values of $T_{2}$, starting the NCG method with a previous solution, and modifying $T_{2}$ in different steps and in opposite directions. The existence of several consecutive unstable branches had already been observed on the CGLE, for instance, in [22] for anomalous 2OD. The profiles of the pulses of those unstable branches are shown in Fig. 6 and all of them exhibit the sharp pulse on top of a larger part that is characteristic of the HAS solutions. Solutions identified as C and D are actually quite similar to the HAS solutions. In the other solutions represented, the larger part of the pulse is seen to be composed of two different regions. This characteristic is more pronounced in pulse F.
@@
MAS and HAS are both stable in a region of parameters, as shown below, producing bistability behaviors. One of these behaviors is hysteresis as may be observed in the results shown on Fig. 6. We solved Eq. (3) using inputs from previous simulations for close values of $T_{2}$. The results labeled as ’Increasing $T_{2}$’ correspond to the output solutions obtained with inputs of lower $T_{2}$ and the ones labeled as ’Decreasing $T_{2}$’ correspond to the output solutions obtained with inputs of higher $T_{2}$. In fact, in the coexistence region the results obtained with different inputs do not coincide. In the direction of decreasing $T_{2}$, the jump between the two type of solutions is abrupt. However, in the other direction, there is a range of $T_{2}$ (105 to 109 fs) for which the simulation did not produce a stationary pulse but instead a pulse that evolved in $Z$ with peak and form variations, showing some similarities with the breathing behaviour that was already reported for dissipative quartic solitons but for negative 4OD [23]. We anticipate that this oscillatory behavior can be explained by the existence of different branches of unstable solutions in this region. Figure 7 shows two kinds of evolution, one converging to a stationary soliton and the other showing the non-stationary evolution referred for the range of $T_{2}$ between 105 to 109 fs.
@@
Figure 6: 6 Bifurcation diagram showing the peak power versus $T_{2}$ with several branches of solutions (including LAS, MAS and HAS). Stable and unstable branches are represented by black and gray lines, respectively. 6 Power profiles for the solutions identified in 6. 6 Hysteresis shown by the peak value of the output pulses obtained by solving Eq. (3) with inputs from higher $T_{2}$ (Decreasing $T_{2}$) and inputs from lower $T_{2}$ (Increasing $T_{2}$). The results shown correspond to $P_{\text{sat}}=80$ W and $g_{0}=1.465$ m-1.
Figure 7: Evolution of peak power and of the full power profile (illustrated by a contour plot) in the case of stationary evolution 7 for $g_{0}=1.465$ m-1 and $T_{2}=100$ fs and non-stationay evolution 7 for $g_{0}=1.465$ m-1 and $T_{2}=107$ fs, both in adimensional variables. The input was the output of a previous simulation for the same $g_{0}$ and a lower $T_{2}$.
@@
Restricting our attention to possible stable solutions, namely, to MAS and HAS solutions, we scanned the $(g_{0},T_{2})$ space to find regions at which those solutions exist and are stable. The regions are those in Figs. 8 and 9 where energy and temporal width are also shown. For increased clarity, the boundaries of the graphs should be explained. The right boundary is at $g_{0}=1.502~{}\text{m}^{-1}$ very close to the limit 1.504 above which the linear loss would become linear gain and the background would become unstable. Above that boundary, even if solitons exist, they would not survive in the unstable background. The upper boundary of MAS, for $T_{2}$ higher than 880 fs and $g_{0}>1.48$, is not an actual boundary but only a limit at which the scanning of the $(g_{0},T_{2})$ space was stopped.
Figure 8: Contour plots of the energy in pJ 8 and pulse width in ps 8 for MAS.
@@
Concerning the region of existence and stability of MAS, it was found to be composed by two distinguishable bands. The first, exists for a $g_{0}$ range between 1.443 and 1.475 m-1, being very limited to $T_{2}$ values around 100 fs. The second band exists in a much broader $T_{2}$ region, with $T_{2}\gtrsim 100$, but is much more limited in $g_{0}$, ranging from 1.475 to 1.5 m-1. Note that the results for $\beta_{4}>0$ presented in [7] were for MAS on the first referred thin band of existence. The lower boundary of the MAS region corresponds to a transition to unstable solutions at which two complex conjugated stability eigenvalues cross the real axis from the stable half plane to the unstable half plane. On the other hand, above the upper boundary, no soliton solutions were found. Figs. 8 and 8 also have contour lines that show that, for lower values of $T_{2}$ the most energetic pulses are also the shortest ones. As noted in [7], the energy increase with $g_{0}$ at a fixed $T_{2}$ is explained by the fact that the former parameter is directly related to energy gain. The energy dependence with $T_{2}$ however is not as straightforward to interpret, as it can be non-monotonous in some cases (assuming constant $g_{0}$). In terms of the width, it can be seen that, for MAS, an increase in $T_{2}$ tends to lead to wider pulses. $T_{2}$ represents the inverse linewidth of the parabolic gain and, therefore, a higher $T_{2}$ could justify larger pulse widths. However, as was shown in Fig. 4, for HAS solutions the width actually decreases with $T_{2}$, showing that there are indeed exceptions to this trend.
@@
Figure 9: Contour plots of base 10 logarithm of the energy in pJ 9 and pulse width in ps 9 for the HAS.
@@
As we referred above, the overlap of regions of stable MAS and HAS is not null, i.e., there are parameter values for which both MAS and HAS solutions exist and are stable. This parameter region of bistability is better observed in Fig. 10, showing it to be small when compared with regions of existence of each separate type of soliton.
@@
Figure 11: Evolution from an LAS solution to a HAS solution (passing through an MAS solution) for $g_{0}=1.47~{}\text{m}^{-1}$ and $T_{2}=100~{}\text{fs}$. 11 Evolution of the peak power in $z$ and 11 power profiles of the three solutions.